A house has 3 windows that are polygons with a total of 13 sides. Two of the windows are the same shape. The third window has one more side than the first two windows. What specific shapes are the windows.

Answer by macston(5194)

.
F=number of sides of first window; S=number of sides of second window=F
T=number of sides of third window=F+1
.
F+S+T=13 Substitute for S and T.
F+F+F+1=13
3F+1=13
3F=12
F=4 The first window has four sides, so it is a quadrilateral.
S=F=4 The second window has four sides, so it is also a quadrilateral.
T=F+1=4+1=5 The third window has five sides, so it is a pentagon.